OpenStudy (anonymous):

Simplify log6 4x +log6 5x + log6 2x. Ca someone please show me how to solve this.

OpenStudy (anonymous):

\[\log _{6} 4x+\log _{6} 5x+\log _{6_{}}2x\]

OpenStudy (anonymous):

utilizing the properties of logs you can write this out as one log with base 6: log_6 (4x*5x*2x)

OpenStudy (anonymous):

\[\log_{a}x+\log_{a}y=\log_{a} xy \]

OpenStudy (anonymous):

that's the property... ^^

OpenStudy (anonymous):

then simplify...

OpenStudy (anonymous):

ohh so its \[\log _{6}40x ^{3}\] ?

OpenStudy (anonymous):

yes